# Maximising Histopathology Segmentation using Minimal Labels via Self-Supervision

### v1

@@ Appendix A Results with Fixed Representations
To evaluate whether self-supervised learning methods are able to learn meaningful representations and generalise to downstream tasks, it is important to use a fixed-feature setting. Therefore, in this setting, the representations learned are used as feature vectors in the downstream tasks of UNet, MDS1, and UDAGAN and the results are provided in Table A, evaluated across different splits of labelled data. Since these self-supervised pre-training methods are not explicitly designed for learning representations well-suited for the tasks of UNet, MDS1, and UDAGAN. Consequently, fixed-feature settings exhibit a significant drop in performance when compared to fine-tuned models (as presented in Table 5), and this is particularly noticeable in the case of HR-CS-CO. This highlights the need for a more effective stain separation methods beyond the classical, matrix decorrelation based approach employed in our study. This is why, during fine-tuning, HR-CS-CO’s representation is able to better adapt to the specific characteristics of the downstream task, and therefore compensate for the limitations caused by the loss of information resulting from the stain separation used during training. Nonetheless, it is essential to acknowledge that even though fixed-feature models experience a decline in performance, they show improved results in comparison to baseline models, especially when employing SimCLR and BYOL as pre-trained models. This improvement is particularly evident when the models are subjected to limited labelled data, such as 1% and 5%. Moreover, when provided with moderate (10%) to fully (100%) labelled data, the fixed-feature models approach the performance levels of baseline models. This highlights the effectiveness of self-supervised learning methods in the context of their capacity to learn meaningful representations.

### v2

@@ B.1 Fixed Representations
To evaluate whether self-supervised learning methods are able to learn meaningful representations and generalise to downstream tasks, it is important to use a fixed-feature setting. Therefore, in this setting, the representations learned are used as feature vectors in the downstream tasks of UNet, MDS1, and UDAGAN and the results are provided in Table B.1, evaluated across different splits of labelled data. Since these self-supervised pre-training methods are not explicitly designed for learning representations well-suited for the tasks of UNet, MDS1, and UDAGAN. Consequently, fixed-feature settings exhibit a significant drop in performance when compared to fine-tuned models (as presented in Table 5), and this is particularly noticeable in the case of HR-CS-CO. This highlights the need for a more effective stain separation methods beyond the classical, matrix decorrelation based approach employed in our study. This is why, during fine-tuning, HR-CS-CO’s representation is able to better adapt to the specific characteristics of the downstream task, and therefore compensate for the limitations caused by the loss of information resulting from the stain separation used during training. Nonetheless, it is essential to acknowledge that even though fixed-feature models experience a decline in performance, they show improved results in comparison to baseline models, especially when employing SimCLR and BYOL as pre-trained models. This improvement is particularly evident when the models are subjected to limited labelled data, such as 1% and 5%. Moreover, when provided with moderate (10%) to fully (100%) labelled data, the fixed-feature models approach the performance levels of baseline models. This highlights the effectiveness of self-supervised learning methods in the context of their capacity to learn meaningful representations.